OpenStudy (anonymous):

simplify logarithm , log base2 8+ log base2 4

OpenStudy (anonymous):

log2(8*4)=log2(32)

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

=log2(2)^5 =5log2(2)=5